# Automatically Labeling \$200B Life-Saving Datasets: A Large Clinical Trial Outcome Benchmark

### v1

@@ 3.1 Agreement with Human Annotations
Table 2: Agreement of aggregated weak labels with TOP Test. $\text{CTO}_{MV}$ denotes Majority Vote, $\text{CTO}_{DP}$ denotes Data Programming [32], and $\text{CTO}_{RF}$ denotes Random Forest. Note that RF actively trained on the training human labels, whereas MV and DP are not.
$\text{CTO}_{MV}$ $\text{CTO}_{DP}$ $\text{CTO}_{RF}$
Phase I II III All I II III All I II III All
F1 0.726 0.689 0.904 0.793 0.870 0.856 0.921 0.884 0.913 0.878 0.941 0.909
PRAUC 0.741 0.688 0.891 0.793 0.819 0.755 0.858 0.807 0.856 0.792 0.894 0.843
ROCAUC 0.751 0.717 0.805 0.775 0.848 0.802 0.743 0.804 0.889 0.838 0.815 0.847
$\kappa$ 0.490 0.430 0.606 0.529 0.700 0.623 0.582 0.646 0.790 0.693 0.710 0.729
We compare our aggregated labels, sourced from various weak labeling methods, against the TOP test set, which is annotated by human experts. The primary goal is to assess the agreement between our labels and the human-annotated labels. To ensure the validity of our analysis, we exclude any trials with a mismatch status between our current dataset and the TOP dataset, as the TOP dataset was annotated using trial data as of 2022. For instance, some trials were labeled as ‘Unknown status’ during that period, whereas in our updated data, they have been marked as ‘completed.’ This also highlights that the trial data changes more often, and annotating them manually each time is not feasible.
Table 2 shows that Data Programming completely beats Majority Vote in terms of agreement with the TOP dataset. Random Forest obtains the highest scores with a Cohen kappa $\kappa$ on all phases of 0.729, indicating substantial agreement [27]. Additionally, the F1 score of 0.909 is much higher than any previous SOTA trial outcome prediction model [11, 41]. Figure 3 also shows that models trained on CTO data performs just as well as those trained on ground truth human annotations.
Figure 3: Comparison of PRAUC performance on the TOP test set with models trained on $\text{CTO}_{RF}$, $\text{CTO}_{DP}$, and ground truth, respectively. We see that the quality of the models using both of our labeling strategies performs similarly to TOP’s training labels, and even occasionally exceeds them.
3.2 Which Labeling Method to Use?
Figure 4: Overview of the phases, statuses and labels of CTO dataset. A) illustrates the distribution of phases within the dataset, highlighting the proportion of each phase. B) presents the distribution of trial statuses, showing how different statuses are represented in the dataset. C) and E) depict the distributions of labels per phase, using random forests and data programming as labeling functions, respectively. D) and F) display the distributions of labels by completion year, also utilizing random forests and data programming as labeling functions. Trials $<$2000 due to their small number.
Determining the better labeling method is subjective and depends on various factors. Figure 4 illustrates the distribution of labels generated by the random forest and data programming labeling functions, showing a significant distribution difference from that of Data Programming (RF predicts almost all trials to be successful). This discrepancy can be attributed to the nature of the random forest labeling function, which relies on a small-scale, human-annotated set that might not capture the overall trial distribution. However, this leads to better agreement with the TOP test set than with data programming.
On the other hand, data programming is an unsupervised approach that generates labels by considering the agreement and relationships between all weak labeling functions across the entire dataset. This approach allows data programming to more accurately learn the trial outcome label distribution. Due to this challenge, we provide both labels from the random forests and the data programming labeling function in our dataset release. Additionally, Figure 4 A and B illustrate the distribution of labels across phases and trial statuses in our proposed CTO dataset. This highlights the sheer quantity of the dataset compared to previous methods.
4 Discussion
We present the first attempt to utilize weak signals to automatically label, a large-scale, reproducible dataset specifically designed for clinical trial outcome prediction. Our dataset and labels facilitate the development of prediction models not only for drug interventions but also for biologics and medical devices, broadening the scope of clinical trial outcome research. Our models consistently perform comparably to baseline models across all phases and metrics when benchmarked against TOP, underscoring the effectiveness of weak supervision and the reliability of our approach. We recognize that automatically created labels will never be a substitute for human ones. However, for data-hungry new ML methods in clinical trial optimization, we assert that this could be a good first step before obtaining human labels, due to our high agreement. In addition, our open-source nature means that any customization to specific tasks can be made quickly and reproducible. Our dataset will be made available at https://github.com/chufangao/CTOD.
Acknowledgments and Disclosure of Funding
@@ 4 Discussion
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Conclusion
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] We are using public data.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Ethics and Broader Impacts
Using a weakly supervised dataset for clinical trial outcome prediction with large language models (LLMs) can potentially decrease the reliability of the model’s predictions if not correctly instantiated with proper labeling functions. Weak supervision may result in incomplete or imprecise labeling, leading to the model learning incorrect associations and missing crucial factors, which can introduce or exacerbate biases. This lack of precise guidance can also cause the model to overfit to noise or incorrect patterns in the training data, reducing its ability to generalize effectively to new, unseen data. In diverse clinical scenarios, one must take care to independently verify model predictions to prevent potentially jeopardizing patient safety with inaccurate predictions. To mitigate these issues, it’s essential to improve the quality of the training data through better labeling techniques, supplementary high-quality data, or advanced methods like semi-supervised or active learning.
Furthermore, we use publicly available data, so the risk of identification is minimized. However, we recognize that LLMs on public data inherently still pose some privacy risks.
Reproducibility
We utilized an AMD EPYC 7513 32-core Processor with 100 GB of RAM to run our experiments. Running Data Programming with less or more powerful systems may impact speed. For us, obtaining the labels took around 25 hours in experiments and prototyping. Additionally, we made use of ChatGPT for writing as well as obtaining trial outcome prediction on the PubMed abstracts. Total cost was around \$200 US dollars.
[…]
1Contents
Appendix B Datasheet for Datasets
@@ A.2    Composition
    See Table 4 and Figure 4.
@@ A.3    Collection process
    We automatically mined each LLM prediction, trial linkage, news headline, and stock price as an overview in Section 2.
@@ A.5    Uses
    Our dataset is designed to promote research primarily in clinical trial outcome prediction. The dataset can also be used for stock price/trend prediction C.3, question answering C.3, etc.

### v2

@@ Results
Table 1: Evaluation of automatically aggregated knowledge base data sources. Agreement of CTO with TOP (on TOP test split). $\text{CTO}_{MV}$ denotes Majority Vote, $\text{CTO}_{DP}$ denotes Data Programming [25], and $\text{CTO}_{RF}$ denotes Random Forest. Note that RF actively trained on the training human labels, whereas MV and DP are not. Bolded denotes best performance.
Phase Aggregation Method F1 $\kappa$
I $\text{CTO}_{MV}$ 0.726 0.490
$\text{CTO}_{DP}$ 0.870 0.700
$\text{CTO}_{RF}$ 0.913 0.790
II $\text{CTO}_{MV}$ 0.689 0.430
$\text{CTO}_{DP}$ 0.856 0.623
$\text{CTO}_{RF}$ 0.878 0.693
III $\text{CTO}_{MV}$ 0.904 0.606
$\text{CTO}_{DP}$ 0.921 0.582
$\text{CTO}_{RF}$ 0.941 0.710
All $\text{CTO}_{MV}$ 0.793 0.529
$\text{CTO}_{DP}$ 0.884 0.646
$\text{CTO}_{RF}$ 0.909 0.729
We evaluate our knowledge base and labeling framework by comparing the agreement between our automatically aggregated labels against the expert-annotated TOP test set. Table 1 summarizes the performance of various label aggregation methods across different trial phases. The label aggregation methods evaluated include Majority Vote (CTOMV), Data Programming (CTODP), and Random Forest-based aggregation (CTORF). Across all trial phases, the Random Forest-based approach consistently outperforms the other methods, achieving the highest scores in both F1 and Cohen’s kappa metrics. CTORF achieves an F1 score of $0.909$ and a Cohen’s kappa ($\kappa$) of $0.729$ when evaluated across all phases, indicating substantial agreement [26]. Also, the phase-specific analysis further underscores the robustness of CTORF. It achieves F1 scores of $0.913$, $0.878$, and $0.941$ in phases I, II, and III, respectively. Additionally, the F1 score of 0.909 is much higher than any previous SOTA trial outcome prediction model [27, 12].
Creating a Manual Curated Trial Outcome Benchmark from CTO
Since automated labeling is never a replacement for humans, we manually annotated over 2,500 challenging cases using our knowledge base to create a high-confidence gold standard set for evaluation. Our process of creating this benchmark is shown in Figure 1D. We follow the same rule-based termination and p-value$<$0.05 as in the automated labeling step to remove the easy cases. For the rest of the unlabeled trials, we utilized 3 clinical trial experts to label the unlabeled trials by assessing related publications, news articles, and other reported outcomes to determine the trial outcome. In this benchmark, we evaluated the labels generated by the CTO automated labeling framework against human labels. The CTO framework demonstrated strong overall performance, achieving an F1 of 97.1. Notably, even in hard and challenging cases (after p-value and status filtering), where accurate labeling is more complex, CTO maintained an F1 of 72.7.
Figure 2: Overview of our manually curated trial outcome benchmark. (A) Bar plot illustrating the top 10 conditions in the benchmark, categorized by MeSH ancestor terms, with associated label counts. (B) Distribution of trial phases and their respective outcome labels. (C) Annual distribution of completed trials in the set from 2020-2024 and their label counts.
Updated Trial Outcome Labels are Highly Valuable
Figure 3: (A) Year-by-year comparison of trial outcome prediction performance for baseline models trained on the TOP dataset, the constantly updated CTO dataset, and our manually curated set, evaluated on trials completed from 2021 to 2024. The TOP dataset is fully utilized for training, whereas for the CTO and manually curated sets, only trials completed prior to each test year are used for training. (B) Top 10 Condition-wise performance of baselines on manual labels, with models trained on data $<$ 2022 and tested on $\geq$ 2022 using TOP, CTO, and manually curated set. (C) Similar to B, Condition-wise performance on the next 10 (Top 11-20 conditions)
Given that we have manual labels for recent trials, we conducted experiments comparing the impact of our constantly updated CTO dataset against the previously publicly available fixed benchmark TOP dataset. We trained standard baseline trial outcome prediction models on the entire TOP dataset (train + valid + test) and assessed their year-by-year performance on trials completed between 2021 and 2024 in our manually curated dataset. For the CTO dataset, the baselines were trained using only trials completed prior to each test year, with training labels generated through our automated labeling framework. Similarly, we trained the baselines on the manually curated dataset using human-labeled outcomes for trials completed before each test year. The phase-wise performance comparison of these baseline models across the different training datasets is presented in Figure 3A.
Baseline models trained using labels from the CTO dataset consistently outperform their counterparts trained on the TOP dataset across all years and trial phases. The CTO dataset significantly enhances model performance compared to the TOP dataset, underscoring the importance of constantly updated labels. This improvement is largely attributable to distribution shifts caused by evolving trial practices, therapeutic areas, and advancements in trial designs over time.
This shows that CTO provides a scalable solution for labeling recent trial data, overcoming the infeasibility and time demands of obtaining new human labels. By achieving high agreement with human labels, CTO enables models to obtain close to human-level performance through leveraging large datasets to adapt to the evolving landscape of clinical trials.
Automated Labels for Condition-Specific Outcome Prediction
Figure 3B and C present the performance of baseline models trained using our human-labeled data across the top 20 conditions in the dataset (3B shows conditions 1-10 and 3C shows conditions 11-20). Here, we train the baselines on trials completed before 2022 over TOP, CTO, and manually curated. The baselines are tested on the manually curated trials completed after 2022. The performance of the baseline models trained on CTO closely follows the models trained on our human labels. This demonstrates that the labels produced by CTO align closely with manual annotations despite being generated independently through an automated process. Moreover, these CTO-based models consistently outperform those trained on the TOP dataset, demonstrating the value of using a more current dataset. Furthermore, CTO facilitates the development of condition-specific clinical trial outcome prediction models by rapidly generating automated labels for condition-specific datasets, significantly reducing the time and effort involved in manual labeling, with CTO-trained baselines matching or even exceeding baselines trained on manually-curated in some cases.
Discussion
Figure 4: Overview of the CTO dataset and labeling functions. (A)Trial distribution by completed year in CTO, which shows a steady increase over the years. (B) Frequency of top 30 conditions in the CTO dataset, categorized by MeSH ancestor terms. (C) Heatmap shows the agreement between the important labeling functions in CTO. (D) Distribution of trial phases and their counts in CTO.
The CTO dataset provides a comprehensive view of clinical trial data, capturing key information across multiple dimensions for around  125K drug and biologics trials. Figure 4A, B, and D provides an overview of the dataset’s key statistics, highlighting its scale and diversity. Figure 4A illustrates the annual distribution of completed clinical trials, and it shows a steady increase over the years, reflecting the growing scale of global clinical research. In Figure 4B, we show the frequency distribution of the top 30 medical conditions within the dataset, where Neoplasms are the most represented condition. The phase distribution of trials in the CTO dataset is shown in Figure 4D. Note that we ensured that our set of manual labels accurately reflects the overall clinical trial distribution as shown in Figure 2A, B and C.
How Well the Labeling Functions Agree?
To assess the agreement between the labeling functions, we calculated the percentage of agreement for trials that fall within the intersections of their coverage, as shown in Figure 4C. The heatmap highlights the agreement scores only for the important labeling functions in the dataset. Each cell in the heatmap represents the agreement score between a pair of labeling functions. It is important to note that each labeling function does not cover all trials with weak labels. Therefore, while calculating the agreement score, we only considered the trials common to both labeling functions. For example, there are no common trials with weak labels between the news headlines and trial status labeling functions (represented as a black cell). This is because the trial status labeling function assigns a value of -1 (indicating an inability to decide the trial outcome) to trials with a ‘completed’ status, while news extraction is limited to ‘completed’ trials to optimize resource usage associated with SerpAPI.
From the heatmap in Figure 2F, several key patterns in labeling function agreement can be observed. Notably, there is a high level of agreement between news headlines and labels derived from p-values, amendments, and serious adverse events. Additionally, we observe a strong agreement between trial linkage and trial status, suggesting that trial linkage effectively identifies failure trials, as most trial status-based labels pertain to trials categorized as terminated,’ withdrawn,’ or ‘suspended.’
Another noteworthy observation is the low agreement between p-values and trial status. This discrepancy arises because most trial status-based failure labels correspond to trials marked as ‘terminated,’ ‘withdrawn,’ or ‘suspended,’ often due to sponsor or business decisions rather than scientific outcomes. Interestingly, many of these trials report statistically significant p-values ($<0.05$) for some of their primary endpoints but are stopped for reasons unrelated to efficacy. For instance, our analysis identified approximately 1,200 trials marked as ‘terminated’ that reported primary outcome p-values below 0.05. Specific examples include NCT04364763, which was terminated due to enrollment feasibility issues, and NCT01374451, which was stopped for failing to meet the primary endpoint of progression-free survival. These cases highlight the complex interplay between trial outcomes and external factors.
Conclusion
We successfully created the first large-scale, publicly available knowledge base of clinical trial outcomes that automatically produces useful outcome labels. We additionally release a set of manually annotated outcome labels using human experts. Since we are the first to publicly release our trial annotation pipeline along with a corresponding knowledge base, we hope to democratize trial outcome prediction, enabling faster, more transparent trial design. Our dataset of recent trials, weak labels, and knowledge base is available at https://chufangao.github.io/CTOD/.
Acknowledgement
@@ A.2    Composition
    Yes, please see Table 3 for a split of training trials with completion date $<2022$ tested on trials with completion date $\geq 2022$.
@@ A.3    Collection process
    We automatically mined each LLM prediction, trial linkage, news headline, and stock price as an overview in Section CTO Overview.
@@ A.5    Uses
    Our dataset is designed to promote research primarily in clinical trial outcome prediction. The dataset can also be used for stock price/trend prediction, question answering, etc.